An airplane accelerates down a runway at 3.75 m/s2 for 32.4s until it finally lifts off the ground. Determine the distance traveled before takeoff.



1
Expert's answer
2020-12-14T12:15:35-0500

"s=at^2/2=3.75\cdot32.4^2/2=1968.3(m)"